25 years ago, a man's age was one-third of what it was 5 years ago. how old is the man now?
Answer by Boreal(15235)

He is x years old
x-25=(1/3)(x-5)
3x-75=x-5, multiplying by 3.
2x=70
x=35 ANSWER
25 years ago, he was 10, and that is 1/3 of what he was 5 years ago, which was 30.